# A Survey of Protocol Fuzzing

### v2

@@ 6.2. Non-Memory-Safety Bug Oracles
Non-Memory-Safety bugs are bugs that are caused by non-memory access reasons and violate some expected properties, such as logical bugs, RFC violations, or performance influential bugs. Non-Memory-Safety bugs are challenging to be identified because they have no uniform observable behavior. Detecting non-memory-safety bugs usually requires the user to define the oracle according to the properties destroyed by the target. Depending on the properties that are checked, these oracles can be roughly divided into four categories, namely incorrect message content & state transitions, inconsistency in transmission, abnormal performance indicators, and differences in execution. We will describe these techniques in detail in the following subsections. It should be noted that although there are various ways to identify possible non-memory-safety bugs, most of these methods can only report suspicious behaviors of the PUT, which still require experts’ manual verification to determine impact and exploitability.
@@ 6.2.1. Incorrect Message Content & State Transitions
Incorrect Message Content checks whether the content of the responses violate some semantic constraints. Incorrect State Transitions verifies whether the state transitions are valid or allowed. In most cases, these rules are extracted from protocol specifications or designed with expert knowledge. These rules can be in different forms, such as canonical state machines (Beurdouche et al., 2015; Somorovsky, 2016), linear-temporal properties (Meng et al., 2021), constraints of response messages (Garbelini et al., 2020, 2022c; Ma et al., 2023), etc. For example, Beurdouche et al. (Beurdouche et al., 2015) manually construct a standard state machine from the specification and then use this state machine as a ground truth to identify deviant behaviors of the PUT. Utilizing this method, a logical bug was identified in a TLS implementation JSSE(Beurdouche et al., 2015). This flaw permitted attackers to bypass all messages pertaining to key exchange and authentication, subsequently enabling them to initiate unencrypted communication. Given a linear-time temporal logic property that a protocol implementation needs to satisfy, LTL-Fuzzer (Meng et al., 2021) leverages directed greybox fuzzing to direct the fuzzing towards specific location that can affect the property, and checks whether the property is held during each execution iteration. Besides, Sweyntooth (Garbelini et al., 2020) and Greyhound (Garbelini et al., 2022c) check whether the received response packet is in the expected type set of the current protocol state. Any mismatched message types are labeled as anomaly. Loki (Ma et al., 2023) extracts rules from the PBFT consensus protocol paper (Castro et al., 1999), which are used as oracles to detect non-memory-safety bugs in blockchain implementations. For example, Loki identified a bug in Hyperledger Fabric (Androulaki et al., 2018) that can be used to confirm illegal transactions.
@@ 6.2.2. Message Inconsistency in Transmission
Some works check whether there are non-memory-safety bugs that can lead to integrity break of the protocol. Specifically, as the correct data transfer is one of the basic properties of TCP protocol, TCP-Fuzz (Zou et al., 2021) designed a data checker on both the sender-side and receiver-side to check the violation of this property. Whenever a message is sent or received, the data checker checks whether the sent message and the received message are identical.
@@ 6.2.3. Abnormal Performance Indicators
Some works aim to find network attack strategies that can affect the performance of the PUT, and these works judge the effectiveness of attack strategies by monitoring whether some performance indicators of PUT are beyond the normal range. For example, to find the amplification DDoS attack strategies in UDP services, AMPFuzz (Krupp et al., 2022) uses the bandwidth amplification factor (BAF) (Rossow, 2014) of each request and response pairs, which is the ratio of the sum of the lengths of all response messages to the length of the attack request, as an indicator to find the message that can maximize the consumption of throughput. TCPWN (Jero et al., 2018) and ABBrate (Peterson et al., 2020) aim to find the attack strategies against the implementations of TCP congestion control that can increase or decrease the congestion window in a model-guided approach. To detect whether the inputs indeed influence the congestion control mechanism, TCPWN obtains the window size from system logs and compares it with an expected baseline.
@@ 6.2.4. Differences in Execution (labeled as “DE” in Table 6 column 6)
Differential testing involves comparing the execution behaviors of different implementations of the same protocol to investigate potential security impacts. This method is scalable due to its independence from code instrumentation. For instance, TCP-Fuzz (Zou et al., 2021) contrasts the outputs of multiple TCP implementations to identify discrepancies. Yang et al. (Yang et al., 2021) employ differential testing to uncover consensus bugs in Ethereum that could lead to fork attacks. They generate a sequence of transactions as inputs and observe the responses of two Ethereum clients, specifically implemented in Golang and Rust. However, a significant challenge in this domain is ascertaining which implementation diverges from the protocol’s expected behavior, and determining whether observed behavioral differences stem from errors or under-specifications in the protocol’s RFC. As such, most works adopting differential testing (Zou et al., 2021; Yang et al., 2021; Bushart and Rossow, 2023) integrate a subsequent manual inspection phase to differentiate actual vulnerabilities from innocuous discrepancies.
To augment the bug-finding efficiency, some studies compare the PUT with an already well-tested or formally verified implementation, referred to as a ‘reference stack’ (Zou et al., 2021). For instance, TCP-Fuzz (Zou et al., 2021) employs classical and extensively tested kernel-level TCP stacks, such as Linux TCP or FreeBSD TCP, as a reference to test newer TCP stacks. In such scenarios, if inconsistencies are reported, it strongly suggests the presence of bugs in the newer protocol implementations. This methodology not only identifies discrepancies but also provides a framework for evaluating the correctness of various protocol implementations.
6.2.5. Timeout and Liveness Checks
@@ 7. Directions of Future Research
So far, we have discussed state-of-the-art protocol fuzzers. In this section, we will answer RQ3 by discussing the research trends and current challenges of fuzzing techniques based on what we have surveyed.
@@ 7.1. Towards Perfect Communication Model Construction
The current methods for constructing communication models are far from perfect, often resulting in either incomplete or inaccurate knowledge acquisition, or requiring extensive manual effort. Specifically, as introduced in Section 4, existing methodologies for constructing communication models can be broadly categorized into top-down and bottom-up approaches. Bottom-up methods are proposed to learn the communication models specific to particular protocol implementations (de Ruiter and Poll, 2015; Zhao et al., 2019; Luo et al., 2019; Fiterau-Brostean et al., 2020; Pham et al., 2020; Wang et al., 2021; Fang et al., 2021; Natella, 2022; Stone et al., 2021; Ba et al., 2022; Qin et al., 2023), rather than the canonical communication model of the protocols themselves. However, for top-down approaches, the majority of existing works still rely heavily on manual processes to construct state machines from protocol specifications (Beurdouche et al., 2015; Jero et al., 2015; Somorovsky, 2016; Kim et al., 2017; Jero et al., 2018; Shi et al., 2019; Garbelini et al., 2020; Ruge et al., 2020; Peterson et al., 2020; Zou et al., 2021; Garbelini et al., 2022c; Park et al., 2022; Romero and Rivas, 2019; He et al., 2022). This manual construction is not only labor-intensive but also prone to errors.
Existing research (Jero et al., 2019; Pacheco et al., 2022) has embarked on exploring the automatic extraction of partial FSMs from protocol specifications using NLP techniques. This exploration has preliminarily validated the feasibility of automating the extraction of protocol communication models. However, this method currently falls short of extracting canonical communication model from protocol specifications due to ambiguities and unspecified behaviors in specifications, thus precluding a complete one-to-one translation between the text and the communication model.
To resolve this, machine learning model based approaches can be explored for better model construction. Considering the recent remarkable progress of LLM (Large Language Model) (Chen et al., 2021), one promising direction is to develop LLM-based solution for more precise model construction. Another possible direction is to combine the other information sources (such as the code of protocol implementations, code commit or comment information during development, program analysis results, etc.) to help understand the content of the specification better.
@@ 7.2. Towards Multi-Dimension Testing Perspectives
Existing research focuses more on changing the content of packets or the order of packet sequences. This approach, while effective to a certain extent, overlooked the fact that protocols have multidimensional testing perspective, e.g., variables such as message latency (Jelassi et al., 2012), cache state (Jung et al., 2001), configurations (Zhang et al., 2023b; Dai et al., 2010), and concurrency level (Katz and Shin, 2006), as highlighted in Section 3.1. These attributes play a crucial role in deciding the behavior of the target system. To effectively test these attributes within protocol implementations, it is necessary to create detailed models that accurately represent each attribute, including message latency, cache state, configuration parameters, and concurrency levels. Additionally, specific oracles and mutators can be designed to evaluate the correctness of the protocol behavior under various scenarios that encompass these multidimensional aspects. This direction is interesting and can help establish a more comprehensive evaluation of the protocol’s resilience and robustness.
@@ 7.3. Fuzzing Characterized Protocol Targets
A significant and under-explored future research direction is the fuzzing of characterized protocol targets. Current research has not comprehensively covered various protocols, especially for those with distinct characteristics and importance. The following three areas are particularly noteworthy:
1. Domain-Specific Protocols. Proprietary domain protocols, such as those used in satellite communication (Sun, 2005), unmanned aerial vehicle (UAV) communication (Gupta et al., 2015), and Robot Operating System (ROS) (Ohkawa et al., 2019), typically possess a high knowledge threshold and a relatively closed nature. These protocols play a crucial role in many infrastructures, making their security research paramount. Presently, fuzzing research for these protocols is relatively scarce, presenting an opportunity for the academic community to improve testing effectiveness and security through the development of new fuzzing techniques and tools.
2. Hardware-Implemented Protocols. Another direction is hardware protocol which designs fuzzers for testing protocols implemented on hardware such as FPGAs (Trimberger and Moore, 2014). These hardware implementations often exhibit different error characteristics compared to those at the software level, necessitating the development of new approaches to more effectively identify and exploit potential vulnerabilities.
3. Multi-Party Protocols. Another possible direction of protocol fuzzing is to support multi-party protocols. In general, protocols have many communication mode, such as peer-to-peer mode (Jero et al., 2018; Walz and Sikora, 2017; Peterson et al., 2020), server-client (master-slave) mode (Vanhoef, 2017; Chen et al., 2018; Fang et al., 2021), and multi-party mode (Sudhodanan et al., 2016). Existing protocol fuzzers focus more on the first two modes by acting as a client/server to test the other (Vanhoef, 2017; Chen et al., 2018; Fang et al., 2021), or playing a role as a peer node to test the PUT (Jero et al., 2018; Walz and Sikora, 2017; Peterson et al., 2020). The multi-party protocols have not been studied. For example, a node can play a role as the computing node, consensus node or management node in a blockchain network (Androulaki et al., 2018), each of which is responsible for a different task. The correct execution of a smart contract protocol requires the cooperation of these roles. How to efficiently test these multi-party protocols is a interesting but challenging question.
@@ 7.4. Combining with Other Vulnerability-Finding Techniques
Beyond fuzzing, there exists a plethora of vulnerability-finding techniques, such as symbolic execution (Sasnauskas et al., 2008; Chau et al., 2017; Asadian et al., 2022; Sun et al., 2018, 2017; Song et al., 2014) and model checking (Musuvathi and Engler, 2004; Jurjens, 2006; Goubault-Larrecq and Parrennes, 2005; Chaki and Datta, 2009; Hoque et al., 2017). While the combination of these techniques with fuzzing has been explored in general contexts (Stephens et al., 2016; Yun et al., 2018), their applications in protocol fuzzing remains relatively under-explored (Song et al., 2014). This presents a promising future research direction, especially considering the fact that the combined approaches still faces the unique testing challenges for complex communications defined in protocols. Intuitively, future research can improve existing vulnerability-finding techniques to better solve protocol-specific challenges. Moreover, many protocols are accompanied by high-quality learning sources, such as detailed specifications. Future research can explore ways to effectively utilize these valuable sources to inform and enhance the combined approaches.
@@ 7.5. Shift-Left Protocol Fuzzing
Though there are certain research efforts focusing on the integration of general-purpose fuzzing techniques into the development cycle – such as with tools like libFuzzer, OSS-Fuzz, and research into fuzzing within CI/CD integration testing (Rangnau et al., 2020) – few studies have specifically dedicated themselves to bridging the gaps between protocol fuzzing and the development process. Protocol fuzzing is distinct from general software fuzzing; it entails rigorously testing the various protocols that allow for communication and data exchange between different software systems and components. Protocol targets generally have a more complex development workflow than that of general software targets. This complexity arises from their need to precisely follow set standards and specifications to ensure interoperability across diverse systems, leading to unique challenges in integration and testing. These challenges necessitate a tailored approach to fuzzing that understands and adapts to the intricacies of protocol development. Therefore, a shift-left approach to protocol fuzzing is needed, which would integrate protocol-specific fuzzing techniques earlier in the software development lifecycle. This can involve the exploration of designing techniques from the developer’s perspective and HCI (Human-Computer Interaction) (Carroll, 1997) techniques can also be considered if necessary. By doing so, it can surface vulnerabilities and issues at an earlier stage where they can be addressed more easily and cost-effectively, ensuring a more robust and secure software ecosystem for protocol implementations.

### v3

@@ 6.2. Non-Memory-Safety Bug Oracles
Non-Memory-Safety bugs are bugs that are caused by non-memory access reasons and violate some expected properties, such as logical bugs, RFC violations, or performance-influential bugs. Non-Memory-Safety bugs are challenging to be identified because they have no uniform observable behavior. Detecting non-memory-safety bugs usually requires the user to define the oracle according to the properties destroyed by the target. Depending on the properties checked, these oracles can be roughly divided into four categories, namely incorrect message content & state transitions, inconsistency in transmission, abnormal performance indicators, and differences in execution. We will describe these techniques in detail in the following subsections. It should be noted that although there are various ways to identify possible non-memory-safety bugs, most of these methods can only report suspicious behaviors of the PUT, which still require experts’ manual verification to determine impact and exploitability.
@@ 6.2.1. Incorrect Message Content & State Transitions
Incorrect Message Content checks whether the content of the responses violates some semantic constraints. Incorrect State Transitions verifies whether the state transitions are valid or allowed. In most cases, these rules are extracted from protocol specifications or designed with expert knowledge. These rules can be in different forms, such as canonical state machines (Beurdouche et al., 2015; Somorovsky, 2016), linear-temporal properties (Meng et al., 2021), constraints of response messages (Garbelini et al., 2020, 2022c; Ma et al., 2023a), etc. For example, Beurdouche et al. (Beurdouche et al., 2015) manually construct a standard state machine from the specification and then use this state machine as a ground truth to identify deviant behaviors of the PUT. Utilizing this method, a logical bug was identified in a TLS implementation JSSE(Beurdouche et al., 2015). This flaw permitted attackers to bypass all messages pertaining to key exchange and authentication, subsequently enabling them to initiate unencrypted communication.
Given a linear-time temporal logic property that a protocol implementation needs to satisfy, LTL-Fuzzer (Meng et al., 2021) leverages directed greybox fuzzing to direct the fuzzing towards specific locations that can affect the property and checks whether the property is held during each execution iteration. In addition, Sweyntooth (Garbelini et al., 2020) and Greyhound (Garbelini et al., 2022c) check whether the received response packet is in the expected type set of the current protocol state. Any mismatched message types are labeled anomaly. Loki (Ma et al., 2023a) extracts rules from the PBFT consensus protocol paper (Castro et al., 1999), which are used as oracles to detect non-memory-safety bugs in blockchain implementations. For example, Loki identified a bug in Hyperledger Fabric (Androulaki et al., 2018) that can be used to confirm illegal transactions.
@@ 6.2.2. Message Inconsistency in Transmission
Some works check whether there are non-memory-safety bugs that can lead to integrity break of the protocol. Specifically, since correct data transfer is one of the basic properties of TCP protocol, TCP-Fuzz (Zou et al., [n. d.]) designed a data checker on both the sender and the receiver side to check the violation of this property. Whenever a message is sent or received, the data checker checks if the sent message and the received message are identical.
@@ 6.2.3. Abnormal Performance Indicators
Some works aim to find network attack strategies that can affect the performance of the PUT, and these works judge the effectiveness of attack strategies by monitoring whether some performance indicators of PUT are beyond the normal range. For example, to find the amplification DDoS attack strategies in UDP services, AMPFuzz (Krupp et al., 2022) uses the bandwidth amplification factor (BAF) (Rossow, 2014) of each pair of requests and responses, which is the ratio of the sum of the lengths of all response messages to the length of the attack request, as an indicator to find the message that can maximize the consumption of throughput. TCPWN (Jero et al., 2018) and ABBrate (Peterson et al., 2020) aim to find attack strategies against implementations of TCP congestion control that can increase or decrease the congestion window in a model-guided approach. To detect whether inputs indeed influence the congestion control mechanism, TCPWN obtains the window size from system logs and compares it with an expected baseline.
@@ 6.2.4. Differences in Execution (labeled as “DE” in Table 5 column 6)
Differential testing involves comparing the execution behaviors of different implementations of the same protocol to investigate potential security impacts. This method is scalable due to its independence from code instrumentation. For example, TCP-Fuzz (Zou et al., [n. d.]) compares the outputs of multiple TCP implementations to identify discrepancies. Yang et al. (Yang et al., 2021) employ differential testing to uncover consensus bugs in Ethereum that could lead to fork attacks. They generate a sequence of transactions as inputs and observe the responses of two Ethereum clients, specifically implemented in Golang and Rust. Similarly, IcyChecker (Ye et al., 2023) identifies blockchain state inconsistency bugs by generating mutated DApp transaction sequences and verifying the consistency of the final states. ParDiff (Zheng et al., 2024) utilizes the bisimulation algorithm to compare the FSMs of different protocol implementations and identifies discrepancies by analyzing the differences in state transition conditions. However, a significant challenge in this domain is ascertaining which implementation diverges from the protocol’s expected behavior, and determining whether observed behavioral differences stem from errors or under-specifications in the protocol’s RFC. As such, most of the work that adopts differential testing (Zou et al., [n. d.]; Yang et al., 2021; Bushart and Rossow, 2023; Kakarla and Beckett, 2023) integrates a subsequent manual inspection phase to differentiate actual vulnerabilities from innocuous discrepancies.
To augment the bug-finding efficiency, some studies compare the PUT with an already well-tested or formally verified implementation, referred to as a ‘reference stack’ (Zou et al., [n. d.]). For example, TCP-Fuzz (Zou et al., [n. d.]) employs classical and extensively tested kernel-level TCP stacks, such as Linux TCP or FreeBSD TCP, as a reference to test newer TCP stacks. In such scenarios, if inconsistencies are reported, it strongly suggests the presence of bugs in the newer protocol implementations. This methodology not only identifies discrepancies, but also provides a framework for evaluating the correctness of various protocol implementations.
@@ 7. Directions of Future Research
So far, we have discussed state-of-the-art protocol fuzzers. In this section, we will address Survey Dimension 3 by discussing the research trends and current challenges of fuzzing techniques based on what we have surveyed.
@@ 7.1. Towards Perfect Communication Model Construction
The current methods for constructing communication models are far from perfect, often resulting in incomplete or inaccurate knowledge acquisition, or requiring extensive manual effort. Specifically, as introduced in Section 4, existing methodologies for constructing communication models can be broadly categorized into top-down and bottom-up approaches. Bottom-up methods are proposed to learn communication models specific to particular protocol implementations (de Ruiter and Poll, 2015; Zhao et al., 2019; Luo et al., 2019; Fiterau-Brostean et al., 2020; Pham et al., 2020; Wang et al., 2021; Fang et al., 2021; Natella, 2022; McMahon Stone et al., 2022; Ba et al., 2022; Qin et al., 2023), rather than the canonical communication model of the protocols themselves. However, for top-down approaches, most existing works still rely heavily on manual processes to construct state machines from protocol specifications (Beurdouche et al., 2015; Jero et al., 2015; Somorovsky, 2016; Kim et al., 2017; Jero et al., 2018; Shi et al., 2019; Garbelini et al., 2020; Ruge et al., [n. d.]; Peterson et al., 2020; Zou et al., [n. d.]; Garbelini et al., 2022c; Park et al., 2022; Romero and Rivas, 2019; He et al., 2022). This manual construction is not only labor intensive, but is also prone to errors.
Existing research (Jero et al., 2019; Pacheco et al., 2022) has begun to explore the automatic extraction of partial FSMs from protocol specifications using NLP techniques. This exploration has preliminarily validated the feasibility of automating the extraction of protocol communication models. However, this method currently falls short of extracting canonical communication models from protocol specifications due to ambiguities and unspecified behaviors in specifications, thus precluding a complete one-to-one translation between the text and the communication model.
To resolve this, machine learning model-based approaches can be explored for better model construction. Considering the recent remarkable progress of LLM (Large Language Model) (Kakarla and Beckett, 2023; Wang et al., 2024; Nikbakht et al., 2024; Zhang et al., 2024), one promising direction is to develop LLM-based solutions for more precise model construction. Another possible direction is to combine the other information sources (such as the code of protocol implementations, code commit or comment information during development, program analysis results, etc.) to help better understand the content of the specification.
@@ 7.2. Towards Multi-Dimension Testing Perspectives
Existing research focuses more on changing the content of packets or the order of packet sequences. This approach, while effective to some extent, overlooked the fact that protocols have multidimensional testing perspectives, e.g., variables such as message latency (Jelassi et al., 2012), cache state (Jung et al., 2001), configurations (Zhang et al., 2023a; Dai et al., 2010), and concurrency level (Katz and Shin, 2006), as highlighted in Section 3.1. These attributes play a crucial role in deciding the behavior of the target system. To effectively test these attributes within protocol implementations, it is necessary to create detailed models that accurately represent each attribute, including message latency, cache state, configuration parameters, and concurrency levels. Additionally, specific oracles and mutators can be designed to assess the correctness of the protocol behavior in various scenarios that encompass these multidimensional aspects. This direction is interesting and can help establish a more comprehensive evaluation of the protocol’s resilience and robustness.
@@ 7.3. Fuzzing Characterized Protocol Targets
A significant and underexplored future research direction is the fuzzing of characterized protocol targets. Current research has not fully covered various protocols, especially for those with distinct characteristics and importance. The following three areas are particularly noteworthy:
1. Domain-Specific Protocols. Proprietary domain protocols, such as those used in satellite communication (Sun, 2014), unmanned aerial vehicle (UAV) communication (Gupta et al., 2015), and Robot Operating System (ROS) (Ohkawa et al., 2019), typically possess a high knowledge threshold and a relatively closed nature. These protocols play a crucial role in many infrastructures, making their security research paramount. Presently, fuzzing research for these protocols is relatively scarce, presenting an opportunity for the academic community to improve testing effectiveness and security through the development of new fuzzing techniques and tools.
2. Hardware-Implemented Protocols. Another direction is to design fuzzers for testing protocols implemented on hardware such as FPGAs (Trimberger and Moore, 2014). These hardware implementations often exhibit different error characteristics compared to those at the software level, necessitating the development of new approaches to more effectively identify and exploit potential vulnerabilities.
3. Multi-Party Protocols. Another possible direction of protocol fuzzing is to support multi-party protocols. In general, protocols have many communication modes, such as peer-to-peer mode (Jero et al., 2018; Walz and Sikora, 2020; Peterson et al., 2020), server-client (master-slave) mode (Vanhoef, 2017; Chen et al., 2018; Fang et al., 2021), and multi-party mode (Sudhodanan et al., 2016). Existing protocol fuzzers focus more on the first two modes by acting as a client/server to test the other (Vanhoef, 2017; Chen et al., 2018; Fang et al., 2021), or playing a role as a peer node to test the PUT (Jero et al., 2018; Walz and Sikora, 2020; Peterson et al., 2020). The multi-party protocols have not been studied. For example, a node can play the role of a computing node, consensus node, or management node in a blockchain network (Androulaki et al., 2018), each of which is responsible for a different task. The correct execution of a smart contract protocol requires the cooperation of these roles. How to efficiently test these multi-party protocols is an interesting but challenging question.
@@ 7.4. Combining with Other Vulnerability-Finding Techniques
Beyond fuzzing, there exists a wide range of vulnerability-finding techniques, such as symbolic execution (Sasnauskas et al., 2008; Chau et al., 2017; Asadian et al., 2022; Sun et al., 2018, 2017; Song et al., 2014) and model checking (Musuvathi and Engler, 2004; Jurjens, 2006; Goubault-Larrecq and Parrennes, 2005; Chaki and Datta, 2009; Hoque et al., 2017). Although the combination of these techniques with fuzzing has been explored in general contexts (Stephens et al., 2016; Yun et al., 2018), their applications in protocol fuzzing remain relatively underexplored (Song et al., 2014). This presents a promising future research direction, especially considering the fact that combined approaches still face unique testing challenges for complex communications defined in protocols. Intuitively, future research can improve existing vulnerability-finding techniques to better solve protocol-specific challenges. In addition, many protocols are accompanied by high-quality learning sources, such as detailed specifications. Future research can explore ways to utilize these valuable sources effectively to inform and enhance combined approaches.
@@ 7.5. Shift-Left Protocol Fuzzing
Although there are certain research efforts focusing on the integration of general-purpose fuzzing techniques into the development cycle – such as with tools like libFuzzer, OSS-Fuzz, and research into fuzzing within CI/CD integration testing (Rangnau et al., 2020) – few studies have specifically dedicated themselves to bridging the gaps between protocol fuzzing and the development process. Protocol fuzzing is distinct from general software fuzzing; it involves rigorously testing the various protocols that allow for communication and data exchange between different software systems and components. Protocol targets generally have a more complex development workflow than general software targets. This complexity arises from their need to precisely follow set standards and specifications to ensure interoperability across diverse systems, leading to unique challenges in integration and testing. These challenges necessitate a tailored approach to fuzzing that understands and adapts to the intricacies of protocol development. Therefore, a shift-left approach to protocol fuzzing is needed, which would integrate protocol-specific fuzzing techniques earlier in the software development lifecycle. This can involve the exploration of designing techniques from the developer’s perspective, and HCI (Human-Computer Interaction) (Carroll, 1997) techniques can also be considered if necessary. By doing so, it can surface vulnerabilities and issues at an earlier stage where they can be addressed more easily and cost-effectively, ensuring a more robust and secure software ecosystem for protocol implementations.